# Collective emission and selective-radiance in atomic clouds and arrays coupled to a microring resonator

### v1

@@ I Introduction
One of the potentials of such atom-nanophotonic interface is the concept of "selective radiance" as was described in Ref. Asenjo-Garcia et al. (2017). These selectively radiant states can be simultaneously superradiant to a desired guided mode and subradiant to undesired or error-inducing modes. Due to the nature of the coupling with the ring resonator, the atoms will superradiantly couple with the resonator, and the decay rate will scale with the atom number Zhou et al. (2024). If the decay into free space can be reduced due to the collective dipole-dipole interactions, the interface becomes more robust against error from spontaneous emission into the vacuum.
@@
This light-matter platform could have potential applications involving quantum memories Gorshkov et al. (2007), quantum simulation Gonzalez-Tudela et al. (2011); Ramos et al. (2014); Pichler et al. (2015); Ramos et al. (2016), quantum networks Stannigel et al. (2012), and applications in quantum computing Paulisch et al. (2016); González-Tudela et al. (2015, 2017). It also holds the potential for investigating photon interactions Roy et al. (2017); Zhang and Mølmer (2019); Sheremet et al. (2023), non-Markovian effects Zhang (2023), and topological models Zhang and Mølmer (2020).
@@
In this paper, we theoretically and computationally model the atom and ring resonator system described in Ref. Zhou et al. (2025) to understand the physics underlying the interactions and explore further possibilities. We give particular emphasis on how the collective dipole-dipole interactions mediated by free space modes can affect the interaction with the microring resonator. We describe the system and the numerical methods used in Sec. II. We simulate the current experiment, which involves a trapped atomic cloud above the ring resonator in Sec. III.2. In Sec. III.3, we explore the potential of an atom array near the ring resonator and consider the effect of disorder that might arise in experiments. This work serves as a companion paper to Ref. Zhou et al. (2025), which describes the experimental aspects and discusses the measured photon decay rates.
II Methods
@@
The experimental system consists of a microring resonator coupled to a bus waveguide, as represented in Fig. 1. The external coupling rate of the resonator to the bus waveguide ($\kappa_{e}$) is comparable to the intrinsic loss rate of the resonator ($\kappa_{i}$), which makes the resonator close to the critical coupling condition Zhou et al. (2023). The microring resonator acts as a whispering gallery mode cavity that can sustain light circulating in the two directions.
Figure 1: Schematic of the experimental system. (a) Cold atoms are trapped above the microring resonator, interacting with a WGM. The WGM is circularly polarized to drive the $\sigma^{+}$ cycling transition |$\ket{g}\rightarrow\ket{e}$. (b,c) Interacting atomic ensemble and organized atom array collectively emit photons into the WGM and the free space modes with photon emission rates $R_{c}$ and $R_{f}$, respectively.
@@
The experiment uses the fundamental Transverse Magnetic-like (TM) WGM. The electric field of the WGM is circularly polarized above the waveguide, and the polarization is locked to the direction of circulation in the resonator. The atoms are polarized to a stretched state and couples to the resonator mode via a cycling transition. As such, the atoms are constrained to emit into the same mode with which it is excited. This configuration results in the atom-light interaction through the waveguide being chiral. We only consider interaction with a fixed direction of circulation, as in Fig. 1(a).
@@
$\frac{d\hat{\rho}}{dt}=-\frac{i}{\hbar}[\hat{H}_{eff},\hat{\rho}]+\mathcal{L}(\hat{\rho}),$ (1)
where $\hbar$ is the reduced Planck constant, $\hat{\rho}$ is the density matrix of the system, $\hat{H}_{eff}$ is the effective Hamiltonian given by Eq. (15) and $\mathcal{L}(\hat{\rho})$ is the Lindblad super-operator which describes the lossy interaction, as defined in Eq. (16).
Since we are restricted to the low-intensity limit, there can be only one excitation in the system at any time. We consider $N$ atoms in the ensemble and the Hilbert space of the atom excitations can be simplified to $N+1$ states. The raising and lowering operators of the $j^{\mathrm{th}}$ atom are represented by $\hat{\sigma}^{+}_{j}$ and $\hat{\sigma}^{-}_{j}$ respectively. The state in which all atoms are in the ground state is represented by $|g\rangle$, and the states where only the $j^{\mathrm{th}}$ atom is excited will be represented by $|e_{j}\rangle=\hat{\sigma}^{+}_{j}|g\rangle$. The single-atom decay rate of the atoms will be denoted by $\Gamma_{0}$ and will serve to set the timescale of the system.
Since there are two types of interactions, we first describe the collective dipole-dipole interactions mediated by the non-guided modes in free space, followed by the effect of coupling the atomic ensemble to the cavity.
II.1 Collective dipole-dipole interaction
Since the separations between adjacent atoms are small and comparable to the wavelength of light, there will be collective dipole-dipole interactions in the absence of the cavity. This can be modeled using the free space dyadic Green’s function $\mathbf{G}(\mathbf{r}_{i},\mathbf{r}_{j},\omega_{0})$, where $\omega_{0}$ is the resonant light frequency and $\mathbf{r}_{i}$, $\mathbf{r}_{j}$ are the position vectors of atom $i$ and $j$.
@@ II.1 Collective dipole-dipole interaction
$\hat{H}_{dd}^{f}=\hbar\sum_{i\neq j}J_{ij}\hat{\sigma}_{i}^{+}\hat{\sigma}_{j}^{-}$ (4)
@@
$\mathcal{L}_{f}(\hat{\rho})=\sum_{i,j}\frac{\Gamma_{ij}}{2}\big{[}2\hat{\sigma}_{i}^{-}\hat{\rho}\hat{\sigma}_{j}^{+}-\hat{\sigma}_{i}^{+}\hat{\sigma}_{j}^{-}\hat{\rho}-\hat{\rho}\hat{\sigma}_{i}^{+}\hat{\sigma}_{j}^{-}\big{]}.$ (5)
@@
The effects of the atoms being in the vicinity of the dielectrics are negligible in this case due to the distance ($z\gtrsim\lambda_{0}/2$, where $\lambda_{0}$ is the resonant wavelength) from the surface (Supplementary of Ref. Zhou et al. (2025)). Therefore, it would suffice to take the simple vacuum Green’s function to make effective qualitative calculations. The interaction matrix $G_{ij}^{f}$ will become
$\begin{split}G^{f}_{ij}=-i\frac{\Gamma_{0}}{2}\bigg{[}h_{0}^{(1)}&(kr_{ij})+\frac{3(\hat{r_{ij}}\cdot\hat{d})(\hat{r_{ij}}\cdot\hat{d}^{*})-1}{2}h_{2}^{(1)}(kr_{ij})\bigg{]},\end{split}$ (7)
@@ II.2 Interaction with the cavity
Once the atomic ensemble is brought to the vicinity of the microring resonator, the atoms can couple with the evanescent field of the cavity and interact. We consider a single mode cavity supporting the $\sigma^{+}$ circularly polarized mode. This cavity interaction can be modeled as a harmonic oscillator with raising (lowering) operators, $\hat{a}^{\dagger}(\hat{a})$. The Hamiltonian of the atom-cavity interaction can be represented as
$\begin{split}\hat{H}_{I}=&\sum_{j}g_{j}(e^{-i\phi_{j}}\hat{a}\hat{\sigma}^{+}_{j}+e^{i\phi_{j}}\hat{a}^{\dagger}\hat{\sigma}^{-}_{j})\\ &-\sum_{j}\Delta_{A}\hat{\sigma}_{j}^{+}\hat{\sigma}_{j}^{-}+\eta(\hat{a}+\hat{a}^{\dagger})-\Delta_{C}\hat{a}^{\dagger}\hat{a},\end{split}$ (8)
@@
In the experiment, single-atom cooperativity plays a more relevant role in describing the strength of the atom-cavity interaction. The single atom cooperativity of atoms is given by,
$C=\frac{4g_{j}^{2}}{(\kappa_{i}+\kappa_{e})\Gamma_{0}},$ (9)
@@
The microring resonator is in the bad-cavity limit, where the cavity dynamics happen at a much faster rate than the dynamics of the atomic internal states ($\kappa_{e},\kappa_{i}\approx 100\Gamma_{0}$). The cavity field can be adiabatically eliminated to more intuitively capture the dynamics of the internal excitations of the atoms. This will also drastically decrease the computational requirements. By enforcing that the cavity reaches a steady state at every time-step of the internal dynamic evolution, we can trace out the cavity modes and obtain a simplified model to study the interactions between the atoms. Further details on the adiabatic eliminations is explained in Appendix A.
@@
$\hat{H}_{dd}^{c}=\hbar\sum_{i\neq j}\mathrm{Re}\{G_{ij}^{c}\}\hat{\sigma}_{i}^{+}\hat{\sigma}_{j}^{-}$ (10)
$\mathcal{L}_{c}(\hat{\rho})=\sum_{i,j}\big{[}2iG_{ij}^{c}\hat{\sigma}_{i}^{-}\hat{\rho}\hat{\sigma}_{j}^{+}+\mathrm{Im}\{G_{ij}^{c}\}(\hat{\sigma}_{i}^{+}\hat{\sigma}_{j}^{-}\hat{\rho}+\hat{\rho}\hat{\sigma}_{i}^{+}\hat{\sigma}_{j}^{-})\big{]},$ (11)
where the interaction matrix through the cavity $G_{ij}^{c}$ will be given by
$\begin{split}G^{c}_{ij}=\frac{g_{i}g_{j}^{*}}{\tilde{\kappa}}e^{-i(\phi_{i}-\phi_{j})},\end{split}$ (12)
and $\tilde{\kappa}=\Delta_{C}+i(\kappa_{i}+\kappa_{e})/2$.
@@
$\hat{H}_{L}=\hbar\sum_{j}\left[-\Delta_{A}\hat{\sigma}_{j}^{+}\hat{\sigma}_{j}^{-}+\frac{\Omega_{j}}{2}\hat{\sigma}_{j}^{+}+\frac{\Omega^{*}_{j}}{2}\hat{\sigma}_{j}^{-}\right],$ (13)
@@
$\Omega_{j}=\frac{g_{j}e^{i\phi_{j}}}{\tilde{\kappa}}\eta.$ (14)
To describe the complete evolution of the system after adiabatic elimination, the effective Hamiltonian and total Lindblad operator in Eq. (1) will be given by
$\hat{H}_{eff}=\hat{H}_{dd}^{f}+\hat{H}_{dd}^{c}+\hat{H}_{L}$ (15)
$\mathcal{L}(\hat{\rho})=\mathcal{L}_{f}(\hat{\rho})+\mathcal{L}_{c}(\hat{\rho}).$ (16)
@@
The density matrix can then be time-evolved using the master equation shown in Eq. (1) with Runge-Kutta methods. The full cavity model and the adiabatically eliminated Lindblad model have both been numerically compared and verified in the single excitation limit. The limitation to adiabatic elimination is that the cavity cannot be in the strong-coupling regime.
@@
An important approximation is that the atoms are stationary within the time scale of the internal excitation decay. This is reasonable since the atoms in the cloud have a mean velocity of 3 cm/s, which corresponds to a couple of nanometers of movement in one excitation lifetime. Hence, to capture the dynamics of the ensemble, we can use a Monte Carlo method to take the average of many different possible positions for the atoms. In each iteration, the positions of the atoms are randomly assigned according to the configuration of the atom cloud.
II.3 Mean field equations
@@
Time evolving the master equation in this particular scenario can lead to a few complications in the calculations. Since the atom cloud is considerably dense, the atoms in the random positions could end up being relatively close to another atom. This will cause the free space Green’s function to increase significantly, beyond the energy scale of the calculation and requiring prohibitively small time-steps. Additionally, since these systems tend to have some states with subradiant character, the time to reach the steady state is much longer than the timescale of typical excitation decay. These issues make it challenging to reach convergence when using more than a few atoms. Hence, we utilize the weak field limit and can derive a procedure to directly calculate the final state expectation values using eigenmode decomposition.
@@
which can be simplified as
@@
where $\vec{\sigma}=\{\langle\hat{\sigma}^{-}_{1}\rangle,...,\langle\hat{\sigma}^{-}_{N}\rangle\}$, $\vec{\Omega}=\{\Omega_{1},...,\Omega_{N}\}$, and $\mathbf{M}=\Delta_{A}\mathbb{1}-\mathbf{G}_{c}-\mathbf{G}_{f}$ is the combined coupling matrix of the two types of interactions.
II.4 Eigenmodes of the excitation
To understand the characteristics of the system, we can study the eigenmodes of the Green’s function of the two different interactions. Apart from having a mismatched wavenumber, the nature of the interaction matrices are different.
For resonant probe and atoms resonant with cavity, $\Delta_{\mathrm{a}}=\Delta_{\mathrm{c}}=0$, $\mathbf{G}_{c}$ is a skew-Hermitian matrix ($\mathbf{G}_{c}^{\dagger}=-\mathbf{G}_{c}$), which results in purely imaginary eigenvalues. There is a single non-zero eigenmode with decay rate $NC_{1}\Gamma_{0}$ with the rest $N-1$ being completely dark subradiant states. In this case, the contribution of $\mathbf{G}_{c}$ to the evolution of atomic coherence $\vec{\sigma}$ will be purely dissipative, which is expected for a resonant atom-cavity interaction. $\mathbf{G}_{f}$, on the other hand, is a complex symmetric matrix ($\mathbf{G}_{f}^{\mathrm{T}}=\mathbf{G}_{f}$) and is both dissipative and dispersive. The combined coupling matrix $\mathbf{M}$ can be diagonalized with N eigenvectors $\vec{v}_{\alpha}$ of complex eigenvalues
$\lambda_{\alpha}=J_{\alpha}-i\frac{\Gamma_{\alpha}}{2}.$ (19)
Here, the $\vec{v}_{\alpha}$ will be the right eigenvectors of the complex matrix $\mathbf{M}$, and further details about the orthogonality relations are explained in Appendix B.
An atomic state can be decomposed using the eigenvectors as
@@ II.4 Eigenmodes of the excitation
where $w_{\alpha}$ is the weight associated with eigenmode $\alpha$.
When driven to the steady state (SS) using a long excitation pulse, the final state $\vec{\sigma}_{\mathrm{SS}}=\sum_{\alpha}w_{\alpha}^{\mathrm{SS}}\vec{v}_{\alpha}$ can be calculated using $\dot{\vec{\sigma}}=0$. In the case of a short pulse, the state can be directly initialized into a timed-Dicke state (TDS) with a small excitation population since the pulse time will be too short to cause mixing into any other states, $\vec{\sigma}_{\mathrm{TDS}}=\sum_{\alpha}w_{\alpha}^{\mathrm{TDS}}\vec{v}_{\alpha}\propto\vec{\Omega}$. The method to calculate $w_{\alpha}^{\mathrm{SS}}$ and $w_{\alpha}^{\mathrm{TDS}}$ is described in Appendix B.

### v2

# Collective emission and selective radiance in atomic clouds and arrays coupled to a microring resonator

@@ I Introduction
One of the potentials of such an atom-nanophotonic interface is the concept of "selective radiance" as was described in Ref. Asenjo-Garcia et al. (2017). These selectively radiant states can be simultaneously superradiant to a desired guided mode and subradiant to undesired or error-inducing modes. Due to the nature of the coupling with the ring resonator, the atoms will superradiantly couple with the resonator, and the decay rate will scale with the atom number Zhou et al. (2024). If the decay into free space can be reduced due to the collective dipole-dipole interactions, the interface becomes more robust against error from spontaneous emission into the vacuum. Such high-wavenumber subradiant states are naturally supported and accessed via the high refractive index ring resonator.
@@
This light-matter platform could also have potential applications involving quantum memories Gorshkov et al. (2007), quantum simulation Gonzalez-Tudela et al. (2011); Ramos et al. (2014); Pichler et al. (2015); Ramos et al. (2016), quantum networks Stannigel et al. (2012), and applications in quantum computing Paulisch et al. (2016); González-Tudela et al. (2015, 2017). It holds the potential for investigating photon interactions Roy et al. (2017); Zhang and Mølmer (2019); Sheremet et al. (2023), non-Markovian effects Zhang (2023), and topological models Zhang and Mølmer (2020).
@@
In this paper, we theoretically and computationally model the atom and ring resonator system described in Ref. Zhou et al. (2025) to understand the physics underlying the interactions and explore further possibilities. We give particular emphasis on how the collective dipole-dipole interactions mediated by free space modes can affect the interaction with the microring resonator. We calculate and compare the decay dynamics of the photons emitted into free space and into the microring resonator.
We describe the system and the numerical methods used in Sec. II and Sec. III, respectively. We simulate the current experiment, which involves a trapped atomic cloud above the ring resonator in Sec. IV.1. In Sec. IV.2, we explore the potential of an atom array near the ring resonator and consider the effect of disorder that might arise in experiments. This work serves as a companion paper to Ref. Zhou et al. (2025), which describes the experimental aspects and discusses the measured photon decay rates.
II The System
@@
The experiment consists of a microring resonator coupled to a bus waveguide, as represented in Fig. 1. The external coupling rate of the resonator to the bus waveguide ($\kappa_{e}$) is comparable to the intrinsic loss rate of the resonator ($\kappa_{i}$), which makes the resonator close to the critical coupling condition Zhou et al. (2023). The microring resonator acts as a whispering gallery mode cavity that can sustain light circulating in the two directions.
[…]
@@
The experiment uses the fundamental Transverse Magnetic-like (TM) WGM. The electric field of the WGM is circularly polarized above the waveguide, and the polarization is locked to the direction of circulation in the resonator. The atoms are polarized to a stretched state and coupled to the resonator mode via a cycling transition. As such, the atoms are constrained to emit into the same mode with which it is excited. This configuration results in the atom-light interaction through the waveguide being chiral. We only consider interaction with a fixed direction of circulation, as in Fig. 1(a).
@@
[…]
$\hat{H}_{{eff}}=\hat{H}_{dd}^{f}+\hat{H}_{dd}^{c}+\hat{H}_{L}^{c}$ (2)
[…]
Since we are restricted to the low-intensity limit, there can be only one excitation in the system at any time. We consider $N$ atoms in the system, and the Hilbert space consists of the collective ground state and the $N$ states in the single excitation manifold. The raising and lowering operators of the $j^{\mathrm{th}}$ atom are represented by $\hat{\sigma}^{+}_{j}$ and $\hat{\sigma}^{-}_{j}$ respectively. The state in which all atoms are in the ground state is represented by $|g\rangle$, and the states where only the $j^{\mathrm{th}}$ atom is excited will be represented by $|e_{j}\rangle=\hat{\sigma}^{+}_{j}|g\rangle$. The single-atom decay rate of the atoms into free space will be denoted by $\Gamma_{0}$ and will serve to set the timescale of the system.
We first describe the collective dipole-dipole interactions mediated by the non-guided modes in free space, followed by the effect of coupling the atomic ensemble to the cavity.
II.1 Free space collective dipole-dipole interaction
The separations between adjacent atoms are small and comparable to the wavelength of light, leading to the emergence of collective dipole-dipole interactions through free space. This can be modeled using the free space dyadic Green’s function $\mathbf{G}(\mathbf{r}_{i},\mathbf{r}_{j},\omega_{0})$, where $\omega_{0}$ is the resonant light frequency and $\mathbf{r}_{i}$, $\mathbf{r}_{j}$ are the position vectors of atom $i$ and $j$.
@@ II.1 Free space collective dipole-dipole interaction
$\hat{H}_{dd}^{f}=\hbar\sum_{i\neq j}J_{ij}\hat{\sigma}_{i}^{+}\hat{\sigma}_{j}^{-},$ (6)
@@
$\mathcal{L}_{f}(\hat{\rho})=\sum_{i,j}\frac{\Gamma_{ij}}{2}\big[2\hat{\sigma}_{j}^{-}\hat{\rho}\hat{\sigma}_{i}^{+}-\hat{\sigma}_{i}^{+}\hat{\sigma}_{j}^{-}\hat{\rho}-\hat{\rho}\hat{\sigma}_{i}^{+}\hat{\sigma}_{j}^{-}\big].$ (7)
@@
The effects of the atoms being in the vicinity of the dielectrics are negligible in this case due to the distance from the surface ($z\gtrsim\lambda_{0}/2$, where $\lambda_{0}$ is the resonant wavelength). More details can be found in the Supplementary Material of Ref. Zhou et al. (2025). Therefore, it would suffice to take the simple vacuum Green’s function to make effective calculations. The interaction matrix $G_{ij}^{f}$ will become
$\begin{split}G^{f}_{ij}=-i\frac{\Gamma_{0}}{2}\bigg[h_{0}^{(1)}&(k_{0}r_{ij})+\frac{3(\hat{r_{ij}}\cdot\hat{d})(\hat{r_{ij}}\cdot\hat{d}^{*})-1}{2}h_{2}^{(1)}(k_{0}r_{ij})\bigg],\end{split}$ (9)
@@ II.2 Interaction with the cavity
[…]
@@
In the experiment, single-atom cooperativity plays a more relevant role in describing the strength of the atom-cavity interaction. The single atom cooperativity of atom $j$ is given by,
$C^{j}=\frac{4g_{j}^{2}}{(\kappa_{i}+\kappa_{e})\Gamma_{0}},$ (11)
@@
The microring resonator is in the bad-cavity limit, where the cavity dynamics happen at a much faster rate than the dynamics of the atomic internal states ($\kappa_{e},\kappa_{i}\approx 100\Gamma_{0}$). The cavity field can be adiabatically eliminated to more intuitively capture the dynamics of the internal excitations of the atoms. This will also drastically decrease the computational requirements. By enforcing that the cavity reaches a steady state at every time step of the internal dynamic evolution, we can trace out the cavity modes and obtain a simplified model to study the interactions between the atoms. Further details on the adiabatic eliminations are explained in Appendix A.
@@
$\hat{H}_{dd}^{c}=\hbar\sum_{i,j}-\mathrm{Re}\bigg\{\frac{g_{i}g_{j}}{\tilde{\kappa}}\bigg\}e^{-i(\phi_{i}-\phi_{j})}\hat{\sigma}_{i}^{+}\hat{\sigma}_{j}^{-}$ (12)
$\begin{split}\mathcal{L}_{c}(\hat{\rho})=\sum_{i,j}&-\mathrm{Im}\bigg\{\frac{g_{i}g_{j}}{\tilde{\kappa}}\bigg\}e^{-i(\phi_{i}-\phi_{j})}\times\\ &\big[2\hat{\sigma}_{j}^{-}\hat{\rho}\hat{\sigma}_{i}^{+}-\hat{\sigma}_{i}^{+}\hat{\sigma}_{j}^{-}\hat{\rho}-\hat{\rho}\hat{\sigma}_{i}^{+}\hat{\sigma}_{j}^{-}\big],\end{split}$ (13)
where $\tilde{\kappa}=\Delta_{C}+i(\kappa_{i}+\kappa_{e})/2$. Similar to the free space Green’s function Eq. (9), we define the interaction matrix through the cavity as
$\begin{split}G^{c}_{ij}=\frac{g_{i}g_{j}}{\tilde{\kappa}}e^{-i(\phi_{i}-\phi_{j})}.\end{split}$ (14)
Excluding the phase dependence $\phi_{i(j)}$ in the ring cavity, the Hamiltonian in Eq. (12) and the dissipation terms in Eq. (13) are dictated by the real and imaginary parts of the interaction matrix in Eq. (14), following the same form as Eqs. (6)–(8).
@@
$\hat{H}_{L}^{c}=\hbar\sum_{j}\left[-\Delta_{A}\hat{\sigma}_{j}^{+}\hat{\sigma}_{j}^{-}+\frac{\Omega_{j}}{2}\hat{\sigma}_{j}^{+}+\frac{\Omega^{*}_{j}}{2}\hat{\sigma}_{j}^{-}\right],$ (15)
@@
$\Omega_{j}=-\frac{g_{j}e^{-i\phi_{j}}}{\tilde{\kappa}}\eta.$ (16)
II.2.1 Evanescent coupling
The atoms couple with the resonator mode through an evanescent coupling. The coupling decreases exponentially with distance from the resonator. The atoms’ average position along the z-axis depends on the trapping potential, allowing control of the average single atom cooperativity $C_{1}$. This provides an opportunity to study the behavior of collective photon emissions as a function of the atom-cavity interaction strength.
The experiment controls the average position of the atoms to be $\sim 400\mathrm{nm}$ away from the surface, which corresponds to $C_{1}\lesssim 0.05$. In the proposed plan to implement an atom array, the trapping methodology can afford the atoms to be much closer to the resonator waveguide, leading to much higher average cooperativities ($C_{1}\sim 10$).
II.3 Comparison of the two interactions
The two types of interactions discussed so far have distinct characteristics that make the system dynamics more interesting and complicated. The eigenmodes and eigenvalues of the corresponding Green’s function matrices can provide an insight into the nature of the interactions. The real part describes the dispersive behavior, and the imaginary part describes the dissipative behavior.
The vacuum mode-mediated interaction between two atoms is symmetric. As a result, the Green’s function of the free space interaction, $\mathbf{G}_{f}$, is a complex symmetric matrix ($\mathbf{G}_{f}^{\mathrm{T}}=\mathbf{G}_{f}$). It typically has non-zero real and imaginary eigenvalues and is hence both dissipative and dispersive. However, the interaction through free-space has a characteristic wavenumber of $k_{0}=2\pi/\lambda_{0}$. This means that excitations which match this wavenumber can couple well with the free space radiation modes. Particularly in ordered arrays, excitations with wavenumber $k>k_{0}$ are beyond the light-cone ($ck>\omega_{0}$) and have suppressed dissipative emission into the vacuum light field.
The resonator-mediated interaction between two atoms is not symmetric, due to the fact that cavity photons circulate around the resonator like running waves. At resonance ($\Delta_{C}=0$), the atom-cavity interaction matrix $\mathbf{G}_{c}$ is a skew-Hermitian matrix ($\mathbf{G}_{c}^{\dagger}=-\mathbf{G}_{c}$), with purely imaginary eigenvalues, making this interaction purely dissipative. There is a single superradiant mode and N1N-1 completely dark subradiant modes. Exciting the atoms through the resonator will always couple to the superradiant mode with decay rate $NC_{1}\Gamma_{0}$ Zhou et al. (2024). The characteristic wavenumber of the atom-cavity interaction will be set by the effective refractive index ($n_{\mathrm{eff}}>1$) of the resonator waveguide, $k_{\mathrm{wg}}=n_{\mathrm{eff}}k_{0}$. This mismatch in wavenumber with the free space interaction causes the superradiant state of the atom-cavity interaction to be subradiant to free space in ordered systems.
III Model for emission dynamics
@@
To understand the dynamics of the system, the density matrix can be time-evolved using the master equation shown in Eq. (1) with Runge-Kutta methods. The full cavity model and the adiabatically eliminated Lindblad model have both been numerically compared and verified in the single excitation limit. The limitation of adiabatic elimination is that the cavity cannot be in the strong-coupling regime.
@@
An important approximation is that the atoms are stationary within the time scale of the internal excitation decay. This is reasonable since the atoms in the cloud have a mean velocity of 3 cm/s, which corresponds to a couple of nanometers of movement in one excitation lifetime. Hence, to capture the dynamics of the ensemble, we can use a Monte Carlo method to take the average of many different possible positions for the atoms. In each iteration, the positions of the atoms are randomly assigned according to the configuration of the atom cloud. This gives rise to a distribution of measured values, as depicted in Fig. 2 (c,d). In all other parts of the manuscript, we focus on the ensemble-averaged values of the relevant observables.
@@
Time evolving the master equation in this particular scenario can lead to a few complications in the calculations. Since the atom cloud is considerably dense, the atoms in the random positions could end up being relatively close to another atom. This will cause the free space Green’s function to increase significantly, beyond the energy scale of the calculation, requiring prohibitively small time-steps. Additionally, since these systems tend to have some states with subradiant character, the time to reach the steady state is much longer than the timescale of typical excitation decay. These issues make it challenging to reach convergence when using more than a few atoms. Hence, we utilize the weak field limit and can derive a procedure to directly calculate the final state expectation values using eigenmode decomposition.
@@
By rewriting the $\langle\hat{\sigma}^{-}_{j}\rangle$ as a vector,
$\vec{\sigma}=\{\langle\hat{\sigma}^{-}_{1}\rangle,...,\langle\hat{\sigma}^{-}_{N}\rangle\},$ (18)
Eq. (17) can be simplified as
@@
where $\vec{\Omega}=\{\Omega_{1},...,\Omega_{N}\}$, and $\mathbf{M}=\Delta_{A}\mathbb{1}-\mathbf{G}_{c}-\mathbf{G}_{f}$ is the combined coupling matrix of the two types of interactions.
This matrix $\mathbf{M}$ can be diagonalized with N eigenvectors $\vec{v}_{\alpha}$ of complex eigenvalues
$\mathbf{M}\vec{v}_{\alpha}=\lambda_{\alpha}\vec{v}_{\alpha},\quad\lambda_{\alpha}=J_{\alpha}-i\frac{\Gamma_{\alpha}}{2}.$ (20)
Here, the $\vec{v}_{\alpha}$ will be the right eigenvectors of the complex matrix $\mathbf{M}$. An atomic state can be decomposed using the eigenvectors as
@@
where $w_{\alpha}$ is the weight associated with eigenmode $\alpha$. The final state after the evolution can be directly solved by separately solving the time evolution of each eigenmode.
III.1 Steady state vs timed-Dicke state
In the experiment in Ref. Zhou et al. (2025), the system is excited using either a short duration or a long duration pulse, creating a single excitation in the system. The collective excited states associated with these situations are considerably different.
In the case of the short pulse, the pulse time (typically around a few nanoseconds) is much shorter than the lifetime of the atomic excitation $(30\mathrm{ns})$. Hence, the pulse can be modelled to be infinitesimally short compared to the timescale of atomic excitation dynamics. The atoms are excited into a state similar to a spin-wave excitation known as the timed-Dicke state (TDS), which is a collective state with a well-defined wave-vector Scully and Svidzinsky (2009). Due to the higher refractive index of the resonator waveguide, the excited TDS will have a wave-vector corresponding to that of the resonator mode $(\mathbf{k}_{\mathrm{wg}})$ denoted by,
$|TDS\rangle=\frac{1}{\sqrt{\mathrm{N}}}\sum_{j}c_{j}e^{i\mathbf{k}_{\mathrm{wg}}\cdot\mathbf{r}_{j}}|e_{j}\rangle,$ (22)
where the coefficient $c_{j}$ is dependent on the individual atom cooperativities $C^{j}$. This will be a superradiant state of the atom-cavity interaction. For decay into free space, the TDS has a higher wavenumber $k_{\mathrm{wg}}$ than $k_{0}$, and is typically associated with subradiance in ordered arrays. For the calculation, the final state can be directly initialized into a timed-Dicke state (TDS) with a small excitation population since the pulse time will be too short to cause mixing into any other states, $\vec{\sigma}_{\mathrm{TDS}}=\sum_{\alpha}w_{\alpha}^{\mathrm{TDS}}\vec{v}_{\alpha}\propto\vec{\Omega}$.
The system can also be driven into the steady state (SS) by using a significantly longer pulse duration ($\sim 200\mathrm{ns}$ in the experiment). Since there is constant driving, the system is usually pumped into excited states exhibiting slower decay. The long excitation facilitates a mixing between different modes, resulting in a range of different wave-vectors, instead of a single dominant wave-vector in the case of the TDS. The detuning also plays an important role in determining which modes are excited in this case. The final state $\vec{\sigma}_{\mathrm{SS}}=\sum_{\alpha}w_{\alpha}^{\mathrm{SS}}\vec{v}_{\alpha}$ can be calculated by assuming an infinitely long excitation until reaching $\dot{\vec{\sigma}}=0$.
The method to calculate $w_{\alpha}^{\mathrm{SS}}$ and $w_{\alpha}^{\mathrm{TDS}}$ and the non-standard orthogonality relations used are described in Appendix B.